OpenStudy (anonymous):

Line AB contains points A (1, 2) and B (−2, 6). The slope of line AB is

OpenStudy (uri):

m=(y2-y1)/(x2-x1) m=6-2/-2-1 4/-3

OpenStudy (anonymous):

Since line AB is passing through two points therefore \[slope -of -AB = \frac{y_1-y_2}{x_1-x_2} = \frac{2-6}{1+2} = \frac{-4}{3} \] @lucy27

OpenStudy (anonymous):

thankyou!!!!!

OpenStudy (anonymous):

The equation of line CD is (y−3) = − 2 (x − 4). What is the slope of a line perpendicular to line CD